A car travels in a straight line along a road. Its distance x from a stop sign is given as a function of time t by the equation π‘₯ = 2.0𝑑2 βˆ’ 0.50𝑑3. Calculate the distance of the car when it achieves its maximum speed in the positive x direction.Β 



1
Expert's answer
2021-05-05T15:01:25-0400


To be asked in question

Maximum speed achieve ix +ive x direction =?

"x=2.0t^2-0.50t^3\rightarrow(1)"

Eqution (1) differenciate with respect to time

"v=\frac{dx}{dt}=4t-1.50t^2\rightarrow(2)"

Eqution (2) differenciate with respect to time

"a=\frac{dv}{dt}=4-3t\rightarrow(3)"

Maximum speed

a=0

"t=\frac{4}{3}"

At "t=\frac{4}{3}" Velocity is maximum then

Displacement

Eqution (1)put values

"x=2\times(\frac{4}{3})^2-0.50\times(\frac{4}{3})^3"

"x=\frac{63}{27}=2.3703meter"